OpenStudy (anonymous):

Input the equation of the given line in standard form. The line including the two points (5, 3), (8, -2)

jimthompson5910 (jim_thompson5910):

were you able to get anywhere with this?

OpenStudy (anonymous):

I kept getting fractions and i feel as if that is wrong

jimthompson5910 (jim_thompson5910):

what did you get for the slope

jimthompson5910 (jim_thompson5910):

the slope is a fraction in this case

jimthompson5910 (jim_thompson5910):

were you able to get the slope?

OpenStudy (anonymous):

(13/2, 1/2) wrong?

jimthompson5910 (jim_thompson5910):

Here's how you find the slope

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (-2-3)/(8-5) m = -5/3 So the slope of the line that passes through the two points is -5/3

jimthompson5910 (jim_thompson5910):

So you can see that the slope is a fraction (and it's perfectly normal to have a fraction for the slope)

jimthompson5910 (jim_thompson5910):

Now turn to y = mx+b form y = mx+b y = (-5/3)x + b ... plug in the given slope 3y = -5x + 3b ... multiply every term by the LCD 3 5x + 3y = 3b ... Add 5x to both sides

jimthompson5910 (jim_thompson5910):

Since this line goes through (5,3), we can plug in x = 5 and y = 3 to find the value of 3b 5x + 3y = 3b 5(5) + 3(3) = 3b 25 + 9 = 3b 34 = 3b 3b = 34 ------------------------------------------------------- So we go from 5x + 3y = 3b to 5x + 3y = 34

jimthompson5910 (jim_thompson5910):

So that's the equation in standard form